OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle?

OpenStudy (anonymous):

use pythagorean theorem

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

a and b are legs and c is the hypotenuse

OpenStudy (anonymous):

so is it 2?

OpenStudy (anonymous):

No, its not 2.

OpenStudy (anonymous):

\[a^{2}+b^{2}=c^{2}\]

OpenStudy (anonymous):

\[11^{2}+b^{2}=12^{2}\]

OpenStudy (anonymous):

\[121+b^{2}=144\]

OpenStudy (anonymous):

\[b^{2}=144-121\]

OpenStudy (anonymous):

\[b^{2}=23\]

OpenStudy (anonymous):

\[\sqrt{b^{2}}=\sqrt{23}\]

OpenStudy (anonymous):

So, \[b=\sqrt{23}\]